OpenStudy (anonymous):

Fan and medal: Assume that the terminal side of an angle of t radians in standard position lies in the given quadrant on the given straight line. Find sin (t), cos (t), tan (t). [Hint: Find a point on the terminal side of the angle.] Quadrant III; line through the origin parallel to 7x - 9y = -2.

OpenStudy (anonymous):

@freckles @DanJS

OpenStudy (danjs):

The point where the line through the origin intersects the unit circle , is the coordinate : \[(x,y) = (\cos(\theta), \sin(\theta))\]

OpenStudy (danjs):

And \[\tan(\theta) = \frac{ y }{ x } = \frac{ \sin(\theta) }{ \cos(\theta) }\]

OpenStudy (king.void.):

This is History! Please close this question and post it in the Math section Thanks!

OpenStudy (danjs):

Oh, i didnt even realize that.

OpenStudy (king.void.):

;)